1.     There are 41 candidates approved to be allotted as the F.D.R.E. council of cabinets. Historic facts of the candidates are shown below, 8 were served as former cabinet members,14 were university presidents, 15 served as members of P.R.H., 2 served in former cabinet and as P.R.H, 4 served as university presidents and members of former cabinet, 6 were served as university presidents and members of P.R.H, 1 served in all the three positions. The find the number of candidates:

a.      Served in none of the three positions?

b.     Only university presidents?

c.      At least one of the three positions?

d.     Exactly two positions?


1
Expert's answer
2022-01-19T14:10:47-0500


a.


"N(C\cup Pr\cup P.R.H)=N(C)+N(Pr)+N(P.R.H)"

"-N(C\cap P.R.H)-N(C\cap Pr)-N(Pr\cap P.R.H)"

"+N(C\cap Pr\cap P.R.H)"

"=8+14+15-2-4-6+1=26"

"N((C\cup Pr\cup P.R.H)^C)=41-N(C\cup Pr\cup P.R.H)"

"=41-26=15"

15 candidates Served in none of the three positions.


b.


"N(only\ Pr)=N(Pr)-N(C\cap Pr)"

"-N(Pr\cap P.R.H)+N(C\cap Pr\cap P.R.H)"

"14-4-6+1=5"

5 candidates served only university presidents.


c. 


"N(C\cup Pr\cup P.R.H)= 26"

26 candidates served at least one of the three positions.   


d.   


"N(exactly\ two\ positiions)"

"=N(C\cap P.R.H)-N(C\cap Pr\cap P.R.H)"

"+N(C\cap Pr)-N(C\cap Pr\cap P.R.H)"

"+N(Pr\cap P.R.H)-N(C\cap Pr\cap P.R.H)"

"=2-1+4-1+6-1=9"

9 candidates served exactly two positions.